An apple juice producer buys apples from a large apple grove that has one variety of apple. The amount of juice squeezed from these apples is approximately normally distributed, with a mean of 4.70 ounces and a standard deviation of 0.40 ounce. Suppose that you select a sample of 25 apples

a) The probability is 70% that the sample mean amount of juice will be contained between what two values symmetrically distributed around the population mean?

b) The probability is 77% that the sample mean amount of juice will be greater than what value?


1
Expert's answer
2021-09-29T00:15:12-0400

Let "X=" the sample mean amount of juice: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=4.70,\sigma=0.40,n=25."


a)

"P(\mu-a<X<\mu+a)=1-2P(X<\mu-a)"

"=1-2P(Z<\dfrac{\mu-a-\mu}{\sigma/\sqrt{n}})=1-2P(Z<\dfrac{-a}{\sigma/\sqrt{n}})"

"=0.70"


"P(Z<\dfrac{-a}{\sigma/\sqrt{n}})=0.15"

"\dfrac{-a}{\sigma/\sqrt{n}}=-1.036433"

"a=1.036433(\dfrac{0.40}{\sqrt{25}})=0.08291464"

"\mu-a=4.70-0.083=4.617"

"\mu+a=4.70+0.083=4.783"

"P(4.617<X<4.783)=0.70"

b)

"P(X>x_2)=1-P(X\leq x_2)"

"=1-P(Z<\dfrac{x_2-\mu}{\sigma/\sqrt{n}})=0.77"

"P(Z<\dfrac{x_2-\mu}{\sigma/\sqrt{n}})=0.23"

"\dfrac{x_2-\mu}{\sigma/\sqrt{n}}=-0.738847"

"x_2=4.70-0.738847(\dfrac{0.40}{\sqrt{25}})=4.641"

"P(X>4.641)=0.77"